EXPERIMENTAL MEASUREMENT OF THE INDUCED DIPOLE MOMENT OF AN ISOLATED MOLECULE IN ITS GROUND AND ELECTRONICALLY EXCITED STATES. INDOLE AND INDOLE-H_{2}O. ^{a}

\underline{CHEOLHWA KANG}, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; TIMOTHY M. KORTER , Department of Chemistry, Syracuse University Syracuse, NY 13244.

Reported here are measurements of the magnitude and orientation of the induced dipole moment that is produced when an indole molecule in its ground S_{0} and electronically excited S_{1} states is polarized by the attachment of a hydrogen bonded water molecule in the gas phase complex indole-H_{2}O. We find the permanent dipole moment values μ_{IW} (S_{0}) = 4.4 and μ_{IW} (S_{1}) = 4.0D, values that are substantially different from calculated values based on vector sums of the dipole moments of the component parts. From this result, we derive the induced dipole moment values μ_{IW}^{∗} (S_{0}) = 0.7 and μ_{IW}^{∗} (S_{1}) = 0.5D. The orientation of the induced moment also is significantly different in the two electronic states. These results are quantitatively reproduced by a purely electrostatic calculation based on ab initio values of multipole moments.

aWork supported by NSF